OpenStudy (narissa):

-4t<_48 @mehek14 with this one i think i have to divide so i believe that answer is t<_-12

OpenStudy (solomonzelman):

\(-4t\le 48\) \(t\ge -12\) Yes, needed to flip the sign when dividing by -4 ....

OpenStudy (narissa):

so i only flip the sign when i divide?

OpenStudy (solomonzelman):

Think about this, if you have \(5\le 8\) And when you divide by -1, you would get \(-5\le -8\) AND THIS IS NOT TRUE! SO the appropriate thing here is to flip the sign, And \(-5\ge -8\), and that would be indeed correct

OpenStudy (solomonzelman):

You flip the sign, when dividng by a nageative

OpenStudy (solomonzelman):

negative **

OpenStudy (narissa):

ohh okay thanks so much

OpenStudy (solomonzelman):

YW